OpenStudy (ijustneedhelp):

what is square root of 7√96m^3

OpenStudy (anonymous):

\[\sqrt{96}=4\sqrt{6}\]

OpenStudy (ijustneedhelp):

but there is a 7 on the outside

OpenStudy (anonymous):

you mean \[\sqrt{7\sqrt{96m^{3}}}\] right?